The compound C4H6N2O6 and hydrogen equals 11.28 grams find the total mass. This is a hypothetical number.
1
Expert's answer
2019-03-18T16:26:40-0400

M(C4H6N2O6)=12*4+6+14*2+16*6 = 48+6+28+96=178 g/mol

for formula : 6g(H) - 178g(C4H6N2O6)

11.28g(H) - x g(C4H6N2O6)

x=178*11.28/6 = 334.64 g (C4H6N2O6)

Answer : 334.64g